# Constraints on Dark Matter with Future MeV Gamma-Ray Telescopes

### v2

@@ III Instrument Details
Here, we list the instrument concepts whose constraints we are forecasting, along with the assumptions we are making about each instrument. In general, unless otherwise noted, we used log interpolation from the data provided in each paper to form our effective area and energy resolution functions. The effective area, energy resolution, and angular resolution as a function of incident photon energy for each instrument is shown in Figure 2. For instruments that did not give an explicit dependence of the energy resolution on incident photon energy, we took the largest value within the range that was given. Each of these instruments has a field-of-view (FOV) between 1 and $2\pi$ sr.
Figure 2: The effective area, energy resolution, and angular resolution as a function of incident photon energy for each of the instruments under consideration. Dotted lines represent the pair-production mode for instruments that have both a Compton mode and a pair-production mode. For GRAMS, we consider the balloon version in our analysis, but there is a proposed upgrade to a satellite version that we also show above using dashed lines for the Compton mode and dot-dashed lines for the pair-production mode. Many of the instruments did not have available a model of the dependence of energy resolution on incident photon energy. For these instruments, we simply used the largest energy resolution given across the entire energy range. The precise definition of angular resolution for each instrument is given in the reference listed for that instrument (see Section III for references).
@@ III.1 GECCO
GECCO (the Galactic Explorer with a Coded Aperture Mask Compton Telescope) combines a Compton telescope with a coded mask telescope for high energy and angular resolution in the range of 0.2 to 8 MeV. An overview of GECCO is given by [22].
@@
e-ASTROGAM (enhanced ASTROGAM) has two modes: a Compton scattering and a pair-production mode. The Compton scattering mode has a low (order 1%) energy resolution, while the pair-production mode has a higher (20-30%) energy resolution. For our bounds, we assumed an energy resolution of 30% ($\Delta E/E$) throughout the pair-production regime. An overview of e-ASTROGAM is given by [20].
@@ III.3 AMEGO
AMEGO (the All-sky Medium Enhanced Gamma-ray Observatory) also has a Compton scattering and a pair-production mode. The Compton scattering mode is divided into two parts: untracked Compton and tracked Compton. This division is responsible for the abrupt jump in the projected decay lifetime constraints. The energy resolution of the Compton mode is given in [23] for each photon energy; for the pair-production mode, we used a conservative assumption of 30% energy resolution ($\Delta E/E$) throughout. An overview of AMEGO is given by [23]. There is a more recent proposal for a similar instrument, AMEGO-X [24]; AMEGO-X has broadly similar effective area and energy resolution to AMEGO, and so we would expect comparable sensitivity, but we show AMEGO results due to the availability of more detailed information on effective area and resolution.
@@ III.4 MAST
MAST (the Massive Argon Space Telescope) is a liquid argon time projection chamber, and covers an energy range of 100 MeV to 1 TeV. An overview of MAST is given by [25].
@@ III.5 AdEPT
AdEPT (the Advanced Energetic Pair Telescope) is a pair-production telescope covering the energy range of 5 to 200 MeV. The effective area at each energy is detailed in [26]; this paper also gives an energy resolution of 30% FWHM at 70 MeV. For our calculations, we assumed that the instrument will have this energy resolution throughout its energy range. An overview of AdEPT is given by [26].
@@ III.6 PANGU
PANGU (the PAir-productioN Gamma-ray Unit) is a pair-production telescope covering the energy range of 10 MeV to 1 GeV. [27] gives an upper bound on the energy resolution of 50% ($\Delta E/E$); we used that value throughout its energy range in our calculations. An overview of PANGU is given by [27].
@@ III.7 GRAMS
GRAMS (the Gamma-Ray and Anti-Matter Survey) uses a liquid argon time-projection chamber and covers the 100 keV to 100 MeV energy range. Its energy resolution has been estimated to be of order 1%, with a slightly higher energy resolution at lower energies [21]. At higher energies, GRAMS has a pair production mode that would provide increased effective area, but worsened energy resolution. For this mode we take the energy resolution to be $50\%$, consistent [28] with the continuum sensitivity analysis of Ref. [21].
An overview of GRAMS is given by [29]. An engineering balloon flight with a small-scale liquid argon time-projection chamber has recently been conducted [30], paving the way for a planned satellite version of GRAMS with detector upgrades. In Figure 2, we plot the effective area and angular resolution for both the balloon (solid for Compton scattering mode, dotted for pair-production mode) and the satellite (dashed for Compton mode, dot-dashed for pair-production mode) versions of GRAMS. The balloon and satellite versions have the same projected energy resolution. In our analysis, we compute the projected sensitivity of the balloon version of GRAMS. The satellite version would improve all sensitivity estimates by a factor of a few due to the increased effective area.
@@ IV.1 Current limits and forecast sensitivity
In Figures 3 and 4, we present the best constraints, across all the instruments considered, for each decay/annihilation mode and target considered.
@@
Figure 3: The best constraints across all the instruments considered for dark matter decay and annihilation into (from top to bottom) two photons, a photon and a neutral pion, and two neutral pions, assuming an observation time of $T_{\mathrm{obs}}=10^{6}\mathrm{\ s}$. These are compared against the current constraints described in the text of Section IV.
@@
Figure 4: The best constraints across all the instruments considered for dark matter decay and annihilation into the final states (from top to bottom) $e^{+}e^{-}$, $\mu^{+}\mu^{-}$, and $\pi^{+}\pi^{-}$, assuming an observation time of $T_{\mathrm{obs}}=10^{6}\mathrm{\ s}$. These are compared against the current constraints described in the text of Section IV.
@@
In Figures 5 and 6, we present our results for each instrument and decay/annihilation mode for observation of the Galactic Center, since this target generally provides the strongest constraints out of all the targets we considered. Results for Draco can be found in Appendix A, and results for M31 can be found simply by rescaling the results for Draco based on the $J$- or $D$-factors of the two targets, since we used the same background model for both targets.
@@
Figure 5: The constraints each instrument, pointed at the Galactic Center, can place on dark matter decay and annihilation into (from top to bottom) two photons, a photon and a neutral pion, and two neutral pions, assuming an observation time of $T_{\mathrm{obs}}=10^{6}\mathrm{\ s}$. These are compared against the current constraints described in Section IV.
@@
Figure 6: The constraints each instrument, pointed at the Galactic Center, can place on dark matter decay and annihilation into the final states (from top to bottom) $e^{+}e^{-}$, $\mu^{+}\mu^{-}$, and $\pi^{+}\pi^{-}$, assuming an observation time of $T_{\mathrm{obs}}=10^{6}\mathrm{\ s}$. These are compared against the current constraints described in the text of Section IV.
@@
For all decay and annihilation channels we considered, current constraints from CMB data from the Planck satellite are given in Ref. [31] ($s$-wave annihilation) and Ref. [32] (decay). For lepton-producing decay and annihilation channels, we use the INTEGRAL constraints from Ref [16], the XMM-Newton constraints from Ref. [33], and the Voyager-1 and AMS-02 constraints given by [34]. Additionally, for decay to electrons and photons, we use the INTEGRAL-SPI (electrons only), COMPTEL, EGRET, and Fermi-LAT given in Ref.[35]. For constraints from INTEGRAL-SPI for decay to photons, we use the data corresponding to the total spectral analysis in 1D parameter space given in Ref. [36]. For the COMPTEL, EGRET, and Fermi-LAT constraints for annihilation to photons as well as for decay or annihilation to a photon and a neutral pion, we use the constraints from Ref. [15].
Where specified, all of the current constraints that we considered are at the 95% confidence limit.
@@ IV.2 Comparison of different experiments
Our forecast constraints from the Galactic Center are stronger than our constraints from M31 / Draco across the board. As one might expect, the Galactic Center seems to be a more promising target for indirect detection of dark matter if the profile is NFW; Draco and M31 provide backup in the event of central densities that are markedly suppressed relative to the NFW profile. Forecast constraints for dark matter decay and annihilation to uncharged particles are generally much stronger than current constraints (except for the current INTEGRAL-SPI constraints given in Ref. [36] in the case of decay to low-energy photons and the current Planck constraints given in Ref. [31] in the case of annihilation to neutral pions, both when considering Draco as a target). Forecast constraints from decay and annihilation to charged particles are weaker than the XMM-Newton constraints given in Ref. [33] (which include inverse Compton emission peaked at lower energies) for decay of dark matter particles with mass higher than about 20 MeV. Additionally, for decay and annihilation to charged particles, forecast constraints from all instruments except MAST are weaker than the Planck constraints given in Ref. [31] for annihilation of dark matter particles with mass higher than about 10 MeV. However, a suppression of annihilation at low velocities, e.g. due to the dominant channel being $p$-wave, would relax the existing Planck constraints relative to those studied in this work.
@@
For the channels including monochromatic photons, MAST offers the most striking sensitivity improvement within the $>100$ MeV energy range, primarily due to its very large projected effective area. Besides MAST, the constraints provided by the other instruments for a given energy range and integration time are all within about an order of magnitude of each other, except for PANGU, which has a particularly large energy resolution. For the instruments that detect lower-energy photons ($<10$ MeV), GECCO has the highest projected sensitivity due to its excellent energy resolution in this band. For the instruments that detect photons with energies between 10 and 100 MeV, AdEPT and GRAMS have the highest projected sensitivity. e-ASTROGAM and AMEGO are also forecast to perform well throughout their energy range. All instruments are forecast to be competitive with or provide improvement over current constraints if pointed at the Galactic Center.
@@
For decay and annihilation channels involving neutral pions, MAST once again offers the strongest forecast constraints due to its large effective area, and e-ASTROGAM and AMEGO are also forecast to perform relatively well. The instruments that are best-suited to detect lower-energy photons (GECCO and GRAMS) are not forecast to provide very strong constraints for channels involving neutral pions, primarily because the photon spectrum from these interactions does not overlap much with their energy range.
@@
For the channels involving annihilation to leptons, MAST is forecast to be competitive with current constraints within its energy range if pointed at the Galactic Center. However, it is not competitive with the XMM-Newton constraints from Ref. [33] for decay. For the channels involving decay or annihilation to electrons, the instruments we consider provide the most promising increase in sensitivity for dark matter masses below about 10 MeV. In this range, particularly in the case of decay or $p$-wave annihilation, GECCO, AMEGO, e-ASTROGAM, and GRAMS are all forecast to offer a sizeable increase in sensitivity to decay or annihilation to electrons when pointed at the Galactic Center (relative to the existing bounds we have considered). None of the instruments considered are particularly well-suited to provide a sensitivity increase for decay or annihilation to charged pions.
@@
We have focused on general final states rather than specific dark matter models in this work. One might ask whether there is a benchmark decay rate or annihilation cross section that these experiments should aim to detect or exclude. We are not aware of such a model-independent target for decay lifetime, and so have preferred to simply show how the forecast sensitivity compares to existing bounds. For $p$-wave annihilation, a possible benchmark would be the cross section that would yield the correct relic density through thermal freezeout. This cross section is somewhat model-dependent, but if freezeout was completely dominated by $p$-wave annihilation and this contribution also dominated the signal today, we would expect the present-day cross section to be of order $\sim 10^{-26}(v_{\text{galactic}}/v_{\text{freezeout}})^{2}$ cm3/s, where $v_{\text{galactic}}$ and $v_{\text{freezeout}}$ denote the typical velocity of dark matter in our Galaxy and at the time of freezeout respectively. Taking $v_{\text{freezeout}}^{2}\sim 0.1$ and $v_{\text{freezeout}}^{2}\sim 10^{-6}$, the target cross section should be in the neighborhood of $\sim 10^{-31}$ cm3/s. This cross section seems very difficult to reach for leptonic final states but may be attainable for photon-rich final states (e.g. involving neutral pions as well as direct annihilation to photons).
@@ IV.3 Systematic uncertainties
We begin our discussion of systematic uncertainty with our signal modeling. Errors in the parameters used to determine the Galactic Center $J$-factors are of order 20% [10], and errors in the parameters used to determine the M31 $J$-factors are of order 10% [11]. We assume throughout that dark matter follows an NFW profile, and there may be larger $J$-factor uncertainties associated with changing the model for the profile. Our signal modeling also assumes that the signal flux comes from one dark matter component which makes up the bulk of dark matter. If this is not the case, our decay / annihilation bounds for any individual dark matter component are too strong by a factor of the fraction of dark matter corresponding to that component (squared in the case of annihilation). For our background modeling, our Fisher analysis takes into account statistical fluctuations and correlations between the background and signal, but does not account for the systematic uncertainties in the background model itself, which may be substantial.
@@
We also remind the reader that as discussed previously, our approach does not use spatial information within the ($10^{\circ}$ radius) ROI, and so very likely underestimates the constraining power of instruments with superior angular resolution. On the other hand, we have neglected instrumental and atmospheric backgrounds, which may degrade the sensitivity estimates in this work if they cannot be reduced below the astrophysical gamma-ray signal.

### v3

@@ III Instrument Details
Here, we list the instrument concepts whose constraints we are forecasting, along with the assumptions we are making about each instrument. In general, unless otherwise noted, we used log interpolation from the data provided in each paper to form our effective area and energy resolution functions. The effective area, energy resolution, and angular resolution as a function of incident photon energy for each instrument is shown in Fig. 2. For instruments that did not give an explicit dependence of the energy resolution on incident photon energy, we took the largest value within the range that was given. Each of these instruments has a field of view between 1 and $2\pi$ sr.
Figure 2: The effective area, energy resolution, and angular resolution as a function of incident photon energy for each of the instruments under consideration. Dotted lines represent the pair-production mode for instruments that have both a Compton mode and a pair-production mode. For GRAMS, we consider the balloon version in our analysis, but there is a proposed upgrade to a satellite version that we also show above using dashed lines for the Compton mode and dot-dashed lines for the pair-production mode. Many of the instruments did not have available a model of the dependence of energy resolution on incident photon energy. For these instruments, we simply used the largest energy resolution given across the entire energy range. The precise definition of angular resolution for each instrument is given in the reference listed for that instrument (see Sec. III for references).
@@ III.1 GECCO
GECCO combines a Compton telescope with a coded mask telescope for high energy and angular resolution in the range of 0.2–8 MeV. The expected energy resolution is $<1^{\circ}$ between 0.5 and 5 MeV, and the expected angular resolution is $\sim 1$ arc min in the $4\times 4\ \mathrm{deg^{2}}$ coded aperture mask data analysis and $4^{\circ}$$8^{\circ}$ in the $60\times 60\ \mathrm{deg^{2}}$ Compton telescope analysis. An overview of GECCO is given by [23].
@@
e-ASTROGAM has two modes: a Compton scattering and a pair-production mode. The Compton scattering mode has a fine (order 1%) energy resolution, while the pair-production mode has a coarser (20%–30%) energy resolution. For our bounds, we assumed an energy resolution of 30% ($\Delta E/E$) throughout the pair-production regime. An overview of e-ASTROGAM is given by [21].
@@ III.3 AMEGO
AMEGO also has a Compton scattering and a pair-production mode. The Compton scattering mode is divided into two parts: untracked Compton and tracked Compton. This division is responsible for the abrupt jump in the projected decay lifetime constraints. The energy resolution of the Compton mode is given in [24] for each photon energy; for the pair-production mode, we used a conservative assumption of 30% energy resolution ($\Delta E/E$) throughout. An overview of AMEGO is given by [24]. There is a more recent proposal for a similar instrument, AMEGO-X [25]; AMEGO-X has broadly similar effective area and energy resolution to AMEGO, and so we would expect comparable sensitivity, but we show AMEGO results due to the availability of more detailed information on effective area and resolution.
@@ III.4 MAST
MAST is a liquid argon time-projection chamber and covers an energy range of 100 MeV–1 TeV. MAST has $\sim 10\%$ energy resolution and $<1^{\circ}$ angular resolution (with much better angular resolution at higher energies) in the energy range that we consider in this work. An overview of MAST is given by [26].
@@ III.5 AdEPT
AdEPT is a pair-production telescope covering the energy range of 5–200 MeV. The effective area at each energy is detailed in [27]; this paper also gives an energy resolution of 30% FWHM at 70 MeV. For our calculations, we assumed that the instrument will have this energy resolution throughout its energy range. An overview of AdEPT is given by [27].
@@ III.6 PANGU
PANGU is a pair-production telescope covering the energy range of 10 MeV–1 GeV. Reference [28] gives an upper bound on the energy resolution of 50% ($\Delta E/E$); we used that value throughout its energy range in our calculations. An overview of PANGU is given by [28].
@@ III.7 GRAMS
GRAMS uses a liquid argon time-projection chamber and covers the 100 keV–100 MeV energy range. Its energy resolution has been estimated to be of order 1%, with a slightly coarser energy resolution at lower energies [22]. At higher energies, GRAMS has a pair-production mode that would provide increased effective area, but worsened energy resolution. For this mode we take the energy resolution to be $50\%$, consistent [29] with the continuum sensitivity analysis of Ref. [22].
An overview of GRAMS is given by [30]. An engineering balloon flight with a small-scale liquid argon time-projection chamber has recently been conducted [31], paving the way for a planned satellite version of GRAMS with detector upgrades. In Fig. 2, we plot the effective area and angular resolution for both the balloon (solid for Compton scattering mode, dotted for pair-production mode) and the satellite (dashed for Compton mode, dot-dashed for pair-production mode) versions of GRAMS. The balloon and satellite versions have the same projected energy resolution. In our analysis, we compute the projected sensitivity of the balloon version of GRAMS. The satellite version would improve all sensitivity estimates by a factor of a few due to the increased effective area.
@@ IV.1 Current limits and forecast sensitivity
In Figs. 3 and 4, we present the best constraints, across all the instruments considered, for each decay/annihilation mode and target considered.
@@
Figure 3: The best constraints across all the instruments considered for dark matter decay and annihilation into (from top to bottom) two photons, a photon and a neutral pion, and two neutral pions, assuming an observation time of $T_{\mathrm{obs}}=10^{6}\mathrm{\ s}$. These are compared against the current constraints described in the text of Sec. IV.
@@
Figure 4: The best constraints across all the instruments considered for dark matter decay and annihilation into the final states (from top to bottom) $e^{+}e^{-}$, $\mu^{+}\mu^{-}$, and $\pi^{+}\pi^{-}$, assuming an observation time of $T_{\mathrm{obs}}=10^{6}\mathrm{\ s}$. These are compared against the current constraints described in the text of Sec. IV.
@@
In Figs. 5 and 6, we present our results for each instrument and decay/annihilation mode for observation of the Galactic Center, since this target generally provides the strongest constraints out of all the targets we considered. Results for Draco can be found in Appendix A, and results for M31 can be found simply by rescaling the results for Draco based on the $J$ or $D$ factors of the two targets, since we used the same background model for both targets.
@@
Figure 5: The constraints each instrument, pointed at the Galactic Center, can place on dark matter decay and annihilation into (from top to bottom) two photons, a photon and a neutral pion, and two neutral pions, assuming an observation time of $T_{\mathrm{obs}}=10^{6}\mathrm{\ s}$. These are compared against the current constraints described in Sec. IV.
@@
Figure 6: The constraints each instrument, pointed at the Galactic Center, can place on dark matter decay and annihilation into the final states (from top to bottom) $e^{+}e^{-}$, $\mu^{+}\mu^{-}$, and $\pi^{+}\pi^{-}$, assuming an observation time of $T_{\mathrm{obs}}=10^{6}\mathrm{\ s}$. These are compared against the current constraints described in the text of Sec. IV.
@@
For all decay and annihilation channels we considered, current constraints from cosmic microwave background data from the Planck satellite are given in Ref. [32] ($s$-wave annihilation) and Ref. [33] (decay). For lepton-producing decay and annihilation channels, we use the INTEGRAL constraints from Ref. [17], the XMM-Newton constraints from Ref. [34], and the Voyager-1 and AMS-02 constraints given by [35]. Additionally, for decay to electrons and photons, we use the INTEGRAL-SPI (electrons only), COMPTEL, EGRET, and Fermi-LAT constraints given in Ref.[36]. For constraints from INTEGRAL-SPI for decay to photons, we use the data corresponding to the total spectral analysis in 1D parameter space given in Ref. [37]. For the COMPTEL, EGRET, and Fermi-LAT constraints for annihilation to photons as well as for decay or annihilation to a photon and a neutral pion, we use the constraints from Ref. [16]. All of the current constraints that we considered are at the 95% confidence limit.
@@ IV.2 Comparison of different experiments
Our forecast constraints from the Galactic Center are stronger than our constraints from M31 / Draco across the board. As one might expect, the Galactic Center seems to be a more promising target for indirect detection of dark matter if the profile is NFW; Draco and M31 provide backup in the event of central densities that are markedly suppressed relative to the NFW profile. Forecast constraints for dark matter decay and annihilation to uncharged particles are generally much stronger than current constraints (except for the current INTEGRAL-SPI constraints given in Ref. [37] in the case of decay to low-energy photons and the current Planck constraints given in Ref. [32] in the case of annihilation to neutral pions, both when considering Draco as a target). Forecast constraints from decay and annihilation to charged particles are weaker than the XMM-Newton constraints given in Ref. [34] (which include inverse Compton emission peaked at lower energies) for decay of dark matter particles with mass higher than about 20 MeV. Additionally, for decay and annihilation to charged particles, forecast constraints from all instruments except MAST are weaker than the Planck constraints given in Ref. [32] for annihilation of dark matter particles with mass higher than about 10 MeV. However, a suppression of annihilation at low velocities, e.g. due to the dominant channel being $p$ wave, would relax the existing Planck constraints relative to those studied in this work.
@@
For the channels including monochromatic photons, MAST offers the most striking sensitivity improvement within the $>100$ MeV energy range, primarily due to its very large projected effective area. With the exception of MAST, the constraints provided by the other instruments for a given energy range and integration time are all within about an order of magnitude of each other, except for PANGU, which has a particularly large energy resolution. For the instruments that detect lower-energy photons ($<10$ MeV), GECCO has the highest projected sensitivity due to its excellent energy resolution in this band. For the instruments that detect photons with energies between 10 and 100 MeV, AdEPT and GRAMS have the highest projected sensitivity. e-ASTROGAM and AMEGO are also forecast to perform well throughout their energy range. All instruments are forecast to be competitive with or provide improvement over current constraints if pointed at the Galactic Center.
@@
For decay and annihilation channels involving neutral pions, MAST once again offers the strongest forecast constraints due to its large effective area, and e-ASTROGAM and AMEGO are also forecast to perform relatively well. The instruments that are best suited to detect lower-energy photons (GECCO and GRAMS) are not forecast to provide very strong constraints for channels involving neutral pions, primarily because the photon spectrum from these interactions does not overlap much with their energy range.
@@
For the channels involving annihilation to leptons, MAST is forecast to be competitive with current constraints within its energy range if pointed at the Galactic Center. However, it is not competitive with the XMM-Newton constraints from Ref. [34] for decay. For the channels involving decay or annihilation to electrons, the instruments we consider provide the most promising increase in sensitivity for dark matter masses below about 10 MeV. In this range, particularly in the case of decay or $p$-wave annihilation, GECCO, AMEGO, e-ASTROGAM, and GRAMS are all forecast to offer a sizable increase in sensitivity to decay or annihilation to electrons when pointed at the Galactic Center (relative to the existing bounds we have considered). None of the instruments considered are particularly well suited to provide a sensitivity increase for decay or annihilation to charged pions.
@@
We have focused on general final states rather than specific dark matter models in this work. One might ask whether there is a benchmark decay rate or annihilation cross section that these experiments should aim to detect or exclude. We are not aware of such a model-independent target for decay lifetime, and so have preferred to simply show how the forecast sensitivity compares to existing bounds. For $p$-wave annihilation, a possible benchmark would be the cross section that would yield the correct relic density through thermal freeze-out. This cross section is somewhat model dependent, but if freeze-out was completely dominated by $p$-wave annihilation and this contribution also dominated the signal today, we would expect the present-day cross section to be of order $\sim 10^{-26}(v_{\text{galactic}}/v_{\text{freeze-out}})^{2}$ cm3/s, where $v_{\text{galactic}}$ and $v_{\text{freeze-out}}$ denote the typical velocity of dark matter in our Galaxy and at the time of freeze-out respectively. Taking $v_{\text{freeze-out}}^{2}\sim 0.1$ and $v_{\text{freeze-out}}^{2}\sim 10^{-6}$, the target cross section should be in the neighborhood of $\sim 10^{-31}$ cm3/s. This cross section seems very difficult to reach for leptonic final states but may be attainable for photon-rich final states (e.g. involving neutral pions as well as direct annihilation to photons).
@@ IV.3 Systematic uncertainties
We begin our discussion of systematic uncertainty with our signal modeling. Errors in the parameters used to determine the Galactic Center $J$ factors are of order 20% [11], errors in the parameters used to determine the M31 $J$ factors are of order 10% [12], and errors in the parameters used to determine the Draco $J$ factors are about 5%–15% [38]. We assume throughout that dark matter follows an NFW profile, and there may be larger $J$ factor uncertainties associated with changing the model for the profile. Our signal modeling also assumes that the signal flux comes from one dark matter component which makes up the bulk of dark matter. If this is not the case, our decay / annihilation bounds for any individual dark matter component are too strong by a factor of the fraction of dark matter corresponding to that component (squared in the case of annihilation). For our background modeling, our Fisher analysis takes into account statistical fluctuations and correlations between the background and signal, but does not account for the systematic uncertainties in the background model itself, which are estimated at 15% [16].
@@
We also remind the reader that as discussed previously, our approach does not use spatial information within the ($10^{\circ}$ radius) ROI, and so very likely underestimates the constraining power of instruments with superior angular resolution. On the other hand, we have neglected instrumental and atmospheric backgrounds, which may degrade the sensitivity estimates in this work if they cannot be reduced below the astrophysical $\gamma$-ray signal.